OpenStudy (dengeki_daisy):

A sphere has center (0,0,0) and a radius of 5. Which of the following points lies on the sphere? Use the distance formula, but adjust it to be used for three variables instead of two. a. (0,-3,4) b. (5,5,-5) c. (2,3,0) d. (1,-1,3)

OpenStudy (dengeki_daisy):

@phi

OpenStudy (phi):

the distance formula between a point (x,y,z) [this is in 3-D] and the point (0,0,0) is \[ \sqrt{x^2+y^2+z^2}\] I would test each of your choices.

OpenStudy (dengeki_daisy):

I'm confused??

OpenStudy (phi):

each of your choices is an (x,y,z) "triplet" square each number in the triplet (i.e. multiply each number by itself) then add up. then find the square root. if you get 5, then you found the correct triplet (correct choice)

OpenStudy (dengeki_daisy):

Is b the answer??

OpenStudy (bobo-i-bo):

Why do you think b is the answer?

OpenStudy (dengeki_daisy):

I squared it

OpenStudy (dengeki_daisy):

@Bobo-i-bo am I right?

OpenStudy (bobo-i-bo):

Please show your working so I can understand what you are doing! And no, you're not right :P

OpenStudy (dengeki_daisy):

\[\sqrt(5^{2}+5^{2}-5^{2}) \]

OpenStudy (dengeki_daisy):

@Bobo-i-bo that's what I did

OpenStudy (bobo-i-bo):

Oh sorry, you are right >.<

OpenStudy (bobo-i-bo):

Yes, you have right answer >.<

OpenStudy (dengeki_daisy):

What did you think it was?

OpenStudy (bobo-i-bo):

Oh no, you did it wrong, sorry you aren't right after all. Your working should be \[\sqrt{5^2+5^2+(-5)^2}\] but I thought you wrote \[\sqrt{5^2+5^2-(5)^2}\] which confused me. You're not right :P

OpenStudy (bobo-i-bo):

You know that we want to find the option which would give us \(5=\sqrt{x^2+y^2+z^2}\) right?

OpenStudy (dengeki_daisy):

OHHHH it's A!!! Yeah I made a mistake I didn't put the -5 in Parentheses, I recalculated it the right way and got 8.66 something.

OpenStudy (dengeki_daisy):

So it's A?? @Bobo-i-bo

OpenStudy (bobo-i-bo):

yes :)

OpenStudy (bobo-i-bo):

you understand why right?

OpenStudy (dengeki_daisy):

@Bobo-i-bo Can you help with one more?? Or do you mind?

OpenStudy (bobo-i-bo):

do it in a new thing :)